# Investigating Primordial Black Hole Accretion through Cosmic Optical Depth

### v1

@@ Abstract
Primordial black holes (PBH) accretion in the late Universe can lead to significant mass growth. A larger mass further accelerates the accretion radiation output for PBHs with initial masses greater than one solar mass, potentially leading to a stringent energy-dumping constraint derived from observations of the cosmic microwave background. The energy injected via PBH accretion is capable of ionizing and heating the intergalactic medium (IGM), ultimately affecting the optical depth of cosmic reionization and the 21-cm signal. This work investigates primordial black hole mass growth using the Bondi-Hoyle accretion model and accounts for additional ionization and heating induced by PBHs. We derive stringent PBH abundance limits using an upper limit on optical depth set by Planck 2018 CMB measurements. We find that accretion growth significantly strengthens late-time observational constraints for primordial black holes with initial masses ranging from several solar masses up to $10^{4}$ solar masses. The PBH fraction of the Universe’s unobserved mass content can be constrained to $f_{\rm PBH,ini}\sim 10^{-2}$ to $10^{-7}$ in this mass range, and when not accounting for mass evolution our constraints can be weakened by up to one order of magnitude. In addition, we show that PBH mass growth will lead to an observable impact on the predicted hydrogen 21-cm brightness temperature.
@@ 1 Introduction
Primordial black holes (PBHs), first proposed by Carr and Hawking in the 1970s [1], can be formed as a result of the collapse of density fluctuations in the early Universe. Depending on the time at which fluctuation modes enter the horizon, as well as later growth [2], primordial black holes may emerge in a wide range of masses [3]. One very intriguing aspect of PBHs is their potential role in the late evolution of the Universe [4]: as massive PBHs barely lose mass via Hawking radiation, they may survive to the present day, thereby contributing to the mass budget that influences galactic dynamics [4]. PBHs can also influence large-scale structure formation in the Universe, affecting matter distribution [5] and many other cosmic evolution features, see Ref [6] for a comprehensive review.
Primordial black holes have been extensively studied as candidates for dark matter (DM), and they are subject to stringent constraints based on astronomical observations. Various methods, including PBH Hawking radiation [7, 8, 9, 10], gravitational lensing [11, 12, 13, 14, 15], cosmic microwave background (CMB) anisotropy [16, 17, 18, 19, 20], gravitational wave [21, 22, 23, 24] and some dynamic methods [25, 26, 27] have been adopted to test PBHs as dark matter in different mass ranges. Projects such as MACHO [13] and EROS [14] have utilized gravitational microlensing to search for PBHs with masses between $10^{-7}$ and $10$ solar masses. Recently, studies using data from the Planck satellite have constrained the abundance of PBHs in the mass range of 10 to $10^{4}$ solar masses [20]. Even at a sub-unity fraction of the invisible mass budget, PBHs are still of strong interest due to other roles in astrophysics and cosmology, e.g. PBHs are predicted to heat the galaxy cores [28], seed supermassive black holes [29], and generate large-scale structure through Poisson fluctuations [5], etc.
The detection of gravitational waves from binary black hole mergers by LIGO and Virgo has revitalized interest in massive PBHs [30, 31, 32, 33, 34]. Both binary PBHs, along with astrophysical populations, may contribute to explaining the observed gravitational wave events. Substantial efforts are underway to investigate the fundamental characteristics of PBHs, including their spin [35], torque [36], and mass distributions [37].
@@ 2.1 The accretion rate of PBHs
satisfies $\chi\geq 2$, namely the PBH Bondi-Hoyle radius is larger than twice the halo radius, one can treat the accretion process of the dark matter halo as a point mass with $M_{\mathrm{h}}$. Otherwise at $\chi<2$ when only a fraction of the halo mass participates in the Bondi-Hoyle accretion, it is necessary to modify the relevant parameters of the point mass accretion model [16] following,
@@
Note for redshifts of interests here, since the DM halo mass is much higher than that of the PBH (see Eq. (2.8)), the velocity dispersion of dark matter is larger than the velocity necessary for escaping PBH. Therefore, the accretion of dark matter to PBHs is negligible and thus not accounted for in this work [16].
@@ 2.3 Mass evolution and boosts to radiation output and accretion rate
In the most extreme case in which $\dot{m}$ consistently takes the upper limit of 10 set by Eddington limit in Eq. (2.17), this equation can be solved analytically and it can be shown that fractional mass evolution $M_{\rm PBH}/M_{{\rm PBH,ini}}$ is dependent only on the age $t$ of the Universe 222 We caution our readers that Eq. (2.22) is only valid for small PBH mass density $\rho_{\rm PBH}$. As we will discuss in detail at next section, if $\rho_{\rm PBH}$ were to be large enough, X-ray from PBH accretion can increase IGM temperature and ionisation, and this may in turn suppress accretion rate (see e.g. Eq. (2.4)) $\dot{m}$ and then Eq. (2.22) will no longer be valid. ,
@@
where $M_{{\rm PBH,ini}}$ is the initial PBH mass, $t_{0}=13.8$ Gyr is the current age of the Universe. When computing mass evolution following Eq. (2.21), we integrate from redshift $z=10^{3}$. We clarify that here the initial values of both PBH mass and abundance are evaluated at $z\geq 10^{3}$, although we note that they may also be extended to $z\sim 350$ because Eq. (2.22) shows that even for the most extreme case, PBH mass will only evolve by about 5% at redshift 350.
Fig. 1 showcases the impacts of accretion on PBH mass and X-ray output. From left to right, we show the evolutions of PBH mass, accretion rate and X-ray luminosity, respectively. We consider three different initial PBH masses $M_{\mathrm{PBH,ini}}$ of 1, 10 and 30 $M_{\odot}$. For convenience, in Fig. 1 we assume that the initial mass density of PBH is a fraction of $10^{-8}$ of that of DM, such that the relevant IGM temperature and ionization remains the same as the $\Lambda{\rm CDM}$ case with no PBH. Due to observational uncertainties in IGM thermal and ionization, we only show results for $z\geq 10$.
@@
To highlight the impact of mass evolution, in the middle and right panels of Fig. 1, we also show results without accounting for mass evolution such that $M_{\rm PBH}=M_{\rm PBH,ini}$ throughout (dotted lines). It can be seen from the red and blue curves that for the same initial mass, as PBH gains more mass, the relevant accretion rate also increases compared with the case of no evolution, i.e. accretion increases not only PBH mass but also accretion rate itself, and this can have a huge impact on accretion X-ray outputs.
In the right panel of Fig. 1, we show accretion X-ray luminosity $L_{\rm X}$ normalized by the initial Eddington luminosity $L_{\rm Edd,ini}\equiv L_{\rm Edd}(M_{\rm PBH,ini})$. Similar to the results for mass evolution and accretion rate, we find that heavier PBHs exhibit higher luminosity. More importantly, once PBH starts to gain mass, the relevant X-ray luminosity shows sharp increase relative to the case of no mass evolution. For $M_{\rm PBH,ini}=30M_{\odot}$, X-ray radiation output can be boosted by a factor of 2000, and for the extreme $\dot{m}=10$ case (blue dotted line), the increase can exceed 4 orders of magnitude.
@@ 3 Impact on the IGM
Fig. 2 showcases the effects of PBH accretion on IGM, from left to right, we show the IGM neutral fraction $x_{\rm HI}\equiv n_{\rm HI}/n_{\rm H}\approx 1-x_{\rm e}$, gas kinetic temperature $T_{\rm K}$ and hydrogen 21cm signals for initial PBH masses between 1 and $10^{4}\ M_{\odot}$. We show results for cases with mass evolution (black solid) and without accounting for mass evolution (red dashed). We assume an initial PBH fraction of $f_{\rm PBH,ini}=10^{-7}$, which is consistent with current observational constraints [6]. From the left panel, it can be seen that at $z=11$, impact of PBH accretion on neutral fraction is negligible when mass evolution is ignored. And when mass growth is accounted for, the boosted X-ray radiation outputs (see Fig. 1) enhances ionization and causes up to 15% decrease in neutral fraction. Since accretion and energy output boost more pronounced for massive PBH, we see that heavier PBHs have larger impact on IGM neutral fraction.
We have shown previously in Fig. 1 that PBH mass growth and boost in X-ray radiation outputs occurs mainly at lower redshifts of $z\leq 30$. Cosmic hydrogen 21-cm signal is particularly suited to explore cosmic ionisation and heating history at this window and can potentially provide valuable insights regarding PBH properties. While the impact on cosmic optical depth primarily derives from the correction on $x_{\rm e}$ history, an additional hydrogen 21-cm constraint can be imposed from the extra heating on the IGM gas temperature. After baryons kinetically decouple from the CMB, neutral hydrogen’s hyperfine state distribution evolves by the balance of multiple spin-flip mechanisms and it is strongly affected by the gas temperature $T_{\rm K}$, especially after the Wouthuysen-Field process takes effect [66]. With extra heating on gas, the 21-cm prediction will change accordingly. The intensity of cosmological 21-cm signal is typically measured by the 21-cm brightness temperature [66]:
@@
Though accounting for stellar X-ray emission and heating is beyond the scope of this paper, we briefly explore the effects of PBH on 21-cm signal assuming saturated Wouthuysen-Field coupling such that $T_{\rm s}=T_{\rm K}$. In the middle and right panels of Fig. 2, we show gas temperature and 21-cm signal, respectively, both plotted at $z=15$. At this redshift, it is observationally viable for astrophysical sources to produce enough Lyman-Alpha photons needed for saturated Wouthuysen-Field coupling, while in the meantime not over-producing X-ray photons to dominate heating history.
Qualitatively speaking, mass growth boosts PBH radiation outputs and leads to enhanced ionization and heating. The ionization effect lowers $x_{\rm HI}$ and can suppress 21-cm signal, however from the left panel we have seen that at $z\geq 11$ the effects of PBH ionization on $x_{\rm HI}$ is relatively small, thus in Fig. 2 the impact of PBH on 21-cm signal is primarily driven by heating. From the middle panel we see that PBH heating increases with the PBH mass, especially for $M_{\mathrm{PBH,ini}}>10M_{\odot}$, and we find that PBH mass evolution can increase gas temperature by a factor of 7 relative to the case without accounting for mass evolution. It is also interesting to note that PBH radiation may even heat $T_{\rm K}$ to be above $T_{\rm CMB}$, potentially flipping the sign of $T_{\rm 21}$ and turn it into an emission signal ($T_{21}>0$). Upcoming measurements can test this in MWA [70], MeerKAT [71] and SKA [72], etc. The full PBH-included $T_{21}$ evolution during EoR requires a more involved simulation, which we will postpone to a follow-up study.
@@ 4 Results
Fig. 3 shows our upper limits on $f_{\rm PBH,ini}$ for $M_{\rm PBH,ini}$ within $[1,10^{4}]M_{\odot}$ range. Our main result is shown in the light blue shaded region, in which we properly account for mass evolution. From our discussions in Sec. 3, it can be inferred that PBH contribution to ionization through deposition term $[{\rm d}E/{\rm d}V{\rm d}t]_{\rm dep,c}$ term is proportional to X-ray emissivity $\epsilon_{\rm X}$, and from Eq. (3.4) and Sec. 2.2 one can easily show that $\epsilon_{\rm X}$ follows the scaling relation,
@@
where $\beta$ equals 2.3 and 0.25 for $\dot{m}\leq 0.02$ and $\dot{m}\geq 0.02$, respectively. $M_{{\rm PBH,ini}}$ scaling of our $f_{{\rm PBH,ini}}$ constraints can be well explained by Eq. (4.3). At smaller $M_{{\rm PBH,ini}}$ when mass growth is negligible ($M_{\rm PBH}/M_{{\rm PBH,ini}}\sim 1$), $\epsilon_{\rm X}$ is dominated by $\dot{m}$ term which is found to increase with PBH mass, thus from the black solid line we see that for $1M_{\odot}\leq M_{{\rm PBH,ini}}<10M_{\odot}$, our $f_{{\rm PBH,ini}}$ limit tightens from $10^{-2}$ to $3\times 10^{-4}$. At higher masses, 44accretion increases both PBH mass ($M_{\rm PBH}/M_{{\rm PBH,ini}}\gg 1$) and accretion rate itself ($\dot{m}$), therefore $f_{{\rm PBH,ini}}$ limit further tightens until $M_{{\rm PBH,ini}}\sim 10^{3}M_{\odot}$. At even higher masses ($M_{{\rm PBH,ini}}\gtrsim 10^{3}M_{\odot}$), accretion rate $\dot{m}$ starts to approach the Eddington limit $\dot{m}=10$. For the most extreme case in which $\dot{m}$ consistently takes $\dot{m}=10$, $M_{\rm PBH,ini}/M_{{\rm PBH,ini}}$ and $\dot{m}$ terms in Eq. (4.3) no longer depends on $M_{{\rm PBH,ini}}$ (see also Eq. (2.22)), thus we see that our constraints gradually flatten at $M_{{\rm PBH,ini}}\gtrsim 10^{3}M_{\odot}$, reaching $f_{{\rm PBH,ini}}\leq 2\times 10^{-7}$ at $M_{{\rm PBH,ini}}=10^{4}M_{\odot}$.
In the dark blue shaded region of Fig. 3, we also show constraints without accounting for mass evolution. For this case Eq. (4.3) reduces to $\epsilon\propto\dot{m}^{\beta}$, since massive PBHs exhibits higher accretion rate $\dot{m}$, our constraint tightens for higher $M_{{\rm PBH,ini}}$ until it gradually flattens when $\dot{m}$ starts to approach the constant upper limit 10. At $M_{\odot}\leq M_{{\rm PBH,ini}}<2M_{\odot}$ the relevant $f_{{\rm PBH,ini}}$ limit is identical to that of with evolution. At higher masses, mass evolution boosts X-ray emissivity, therefore we see that when ignoring mass growth, $f_{{\rm PBH,ini}}$ constraints can be weakened by up to one order of magnitude.

### v2

@@ Abstract
Primordial black holes (PBH) accretion in the late Universe can lead to significant mass growth. A larger mass further accelerates the accretion radiation output for PBHs with initial masses greater than one solar mass, potentially leading to a stringent energy-dumping constraint derived from observations of the cosmic microwave background. The energy injected via PBH accretion is capable of ionizing and heating the intergalactic medium (IGM), ultimately affecting the optical depth of cosmic reionization and the 21-cm signal. This work investigates primordial black hole mass growth using the Bondi-Hoyle accretion model and accounts for additional ionization and heating induced by PBHs. We derive stringent PBH abundance limits using an upper limit on optical depth set by Planck 2018 CMB measurements. We find that accretion growth significantly strengthens late-time observational constraints for primordial black holes with initial masses ranging from several solar masses up to $10^{4}$ solar masses. The PBH fraction of the Universe’s unobserved mass content can be constrained to $f_{\rm PBH,ini}\sim 10^{-2}$ to $10^{-7}$ in this mass range, and when accounting for mass evolution our constraints can be strengthened by up to one order of magnitude. In addition, we show that PBH mass growth will lead to an observable impact on the predicted hydrogen 21-cm brightness temperature.
@@ 1 Introduction
Primordial black holes (PBHs), first proposed by Carr and Hawking in the 1970s [1], can be formed as a result of the collapse of density fluctuations in the early Universe. Depending on the time at which fluctuation modes enter the horizon, as well as later growth [2], primordial black holes may emerge in a wide range of masses [3, 4]. One very intriguing aspect of PBHs is their potential role in the late evolution of the Universe [5]: as massive PBHs barely lose mass via Hawking radiation, they may survive to the present day, thereby contributing to the mass budget that influences galactic dynamics [5]. PBHs can also influence large-scale structure formation in the Universe, affecting matter distribution [6] and many other cosmic evolution features, see Ref [7] for a comprehensive review.
Primordial black holes have been extensively studied as candidates for dark matter (DM), and they are subject to stringent constraints based on astronomical observations. Various methods, including PBH Hawking radiation [8, 9, 10, 11], gravitational lensing [12, 13, 14, 15, 16, 17], cosmic microwave background (CMB) anisotropy [18, 19, 20, 21, 22, 23, 24], gravitational wave [25, 26, 27, 28, 29] and some dynamic methods [30, 31, 32] have been adopted to test PBHs as dark matter in different mass ranges. Projects such as MACHO [14] and EROS [15] have utilized gravitational microlensing to search for PBHs with masses between $10^{-7}$ and $10$ solar masses. Recently, studies using data from the Planck satellite have constrained the abundance of PBHs in the mass range of 10 to $10^{4}$ solar masses [22]. Even at a sub-unity fraction of the invisible mass budget, PBHs are still of strong interest due to other roles in astrophysics and cosmology, e.g. PBHs are predicted to heat the galaxy cores [33], seed supermassive black holes [34] and galaxies [35], and generate large-scale structure through Poisson fluctuations [6], etc.
The detection of gravitational waves from binary black hole mergers by LIGO and Virgo has revitalized interest in massive PBHs [36, 37, 38, 39, 40, 41, 42, 43]. Both binary PBHs, along with astrophysical populations, may contribute to explaining the observed gravitational wave events. Substantial efforts are underway to investigate the fundamental characteristics of PBHs, including their spin [44], torque [45], and mass distributions [46].
@@ 2.1 The accretion rate of PBHs
satisfies $\chi\geq 2$, namely, the PBH Bondi-Hoyle radius is larger than twice the halo radius, and one can treat the accretion process of the dark matter halo as a point mass with $M_{\mathrm{h}}$. Otherwise, at $\chi<2$ when only a fraction of the halo mass participates in the Bondi-Hoyle accretion, it is necessary to modify the relevant parameters of the point mass accretion model [18] following,
@@
Note that the accretion of dark matter to PBHs is negligible [18] and thus not accounted for in this work. The velocity dispersion of dark matter is typically much higher than the escape velocity required for dark matter to be gravitationally bound to the PBH, this dynamical imbalance implies that the random thermal motions of DM particles dominate over the PBH’s gravitational pull, rendering direct accretion inefficient. Furthermore, the density profile of the DM halo extends over scales orders of magnitude larger than the Schwarzschild radius of the PBH. Such vast spatial separation introduces significant angular momentum barriers, and for DM to spiral into the PBH they must first lose nearly all their angular momentum, which is highly improbable given the lack of efficient dissipative mechanisms for DM.
@@ 2.3 Mass evolution and boosts to radiation output and accretion rate
For sufficiently large accretion rate and PBH abundance, PBH accretion may consume a substantial fraction of baryonic matter and this can in turn suppress the accretion rate. For the regime of interest in this work, we have checked that the fractional loss in baryonic gas is negligible (at up to $10^{-2}$ level), thus we ignore this effect in our subsequent analysis.
In the most extreme case in which $\dot{m}$ consistently takes the upper limit of 10 set by Eddington limit in Eq. (2.17), Eq. (2.21) can be solved analytically and it can be shown that fractional mass evolution $M_{\rm PBH}/M_{{\rm PBH,ini}}$ is dependent only on the age $t$ of the Universe 222 We caution our readers that Eq. (2.22) is only valid for small PBH mass density $\rho_{\rm PBH}$. As we will discuss in detail at next section, if $\rho_{\rm PBH}$ were to be large enough, X-ray from PBH accretion can increase IGM temperature and ionisation, and this may in turn suppress accretion rate (see e.g. Eq. (2.4)) $\dot{m}$ and then Eq. (2.22) will no longer be valid. ,
@@
where $t\in[0,t_{0}]$, corresponding to redshift range of $[0,\infty]$. $M_{{\rm PBH,ini}}$ is the initial PBH mass, $t_{0}=13.8$ Gyr is the current age of the Universe. When computing mass evolution for realistic accretion scenarios detailed in Sec. 2.1, we integrate Eq. (2.21) from redshift $z=10^{3}$. We clarify that here the initial values of both PBH mass and abundance are evaluated at $z\geq 10^{3}$, although we note that they may also be extended to $z\sim 350$ because Eq. (2.22) shows that even for the most extreme case, PBH mass will only evolve by about 5% at redshift 350.
Figure 1 showcases the impacts of accretion on PBH mass and X-ray output. From left to right, we show the evolutions of PBH mass, accretion rate and X-ray luminosity, respectively. We consider three different initial PBH masses $M_{\mathrm{PBH,ini}}$ of 1, 10 and 30 $M_{\odot}$. For convenience, in Figure 1 we assume that the initial mass density of PBH is a fraction of $10^{-8}$ of that of DM, such that the relevant IGM temperature and ionization remains the same as the $\Lambda{\rm CDM}$ case with no PBH. Due to observational uncertainties in IGM thermal and ionization, we only show results for $z\geq 10$.
@@
To highlight the impact of mass evolution, in the middle and right panels of Figure 1, we also show results without accounting for mass evolution such that $M_{\rm PBH}=M_{\rm PBH,ini}$ throughout (dotted lines). It can be seen from the red and blue curves that for the same initial mass, as PBH gains more mass, the relevant accretion rate also increases compared with the case of no evolution, i.e. accretion increases not only PBH mass but also accretion rate itself, and this can have a huge impact on accretion X-ray outputs.
In the right panel of Figure 1, we show accretion X-ray luminosity $L_{\rm X}$ normalized by the initial Eddington luminosity $L_{\rm Edd,ini}\equiv L_{\rm Edd}(M_{\rm PBH,ini})$. Similar to the results for mass evolution and accretion rate, we find that heavier PBHs exhibit higher luminosity. More importantly, once PBH starts to gain mass, the relevant X-ray luminosity shows sharp increase relative to the case of no mass evolution. For $M_{\rm PBH,ini}=30M_{\odot}$, X-ray radiation output can be boosted by a factor of 2000, and for the extreme $\dot{m}=10$ case (blue dotted line), the increase can exceed 4 orders of magnitude.
@@ 3 Impact on the IGM
Figure 2 showcases the effects of PBH accretion on IGM, from left to right, we show the IGM neutral fraction $x_{\rm HI}\equiv n_{\rm HI}/n_{\rm H}\approx 1-x_{\rm e}$, gas kinetic temperature $T_{\rm K}$ and hydrogen 21cm signals for initial PBH masses between 1 and $10^{4}\ M_{\odot}$. We show results for cases with mass evolution (black solid) and without accounting for mass evolution (red dashed). We assume an initial PBH fraction of $f_{\rm PBH,ini}=10^{-7}$, which is consistent with current observational constraints [7]. From the left panel, it can be seen that at $z=11$, impact of PBH accretion on neutral fraction is negligible when mass evolution is ignored. And when mass growth is accounted for, the boosted X-ray radiation outputs (see Figure 1) enhances ionization and causes up to 15% decrease in neutral fraction. Since accretion and energy output boost more pronounced for massive PBH, we see that heavier PBHs have larger impact on IGM neutral fraction.
We have shown previously in Figure 1 that PBH mass growth and boost in X-ray radiation outputs occurs mainly at lower redshifts of $z\leq 30$. Cosmic hydrogen 21-cm signal is particularly suited to explore cosmic ionisation and heating history at this window and can potentially provide valuable insights regarding PBH properties. While the impact on cosmic optical depth primarily derives from the correction on $x_{\rm e}$ history, an additional hydrogen 21-cm constraint can be imposed from the extra heating on the IGM gas temperature. After baryons kinetically decouple from the CMB, neutral hydrogen’s hyperfine state distribution evolves by the balance of multiple spin-flip mechanisms and it is strongly affected by the gas temperature $T_{\rm K}$, especially after the Wouthuysen-Field process takes effect [73]. With extra heating on gas, the 21-cm prediction will change accordingly. The intensity of cosmological 21-cm signal is typically measured by the 21-cm brightness temperature [73]:
@@
Though accounting for stellar X-ray emission and heating is beyond the scope of this paper, we briefly explore the effects of PBH on 21-cm signal assuming saturated Wouthuysen-Field coupling such that $T_{\rm s}=T_{\rm K}$. In the middle and right panels of Figure 2, we show gas temperature and 21-cm signal, respectively, both plotted at $z=15$. At this redshift, it is observationally viable for astrophysical sources to produce enough Lyman-Alpha photons needed for saturated Wouthuysen-Field coupling, while in the meantime not over-producing X-ray photons to dominate heating history.
Qualitatively speaking, mass growth boosts PBH radiation outputs and leads to enhanced ionization and heating. The ionization effect lowers $x_{\rm HI}$ and can suppress 21-cm signal, however from the left panel we have seen that at $z\geq 11$ the effects of PBH ionization on $x_{\rm HI}$ is relatively small, thus in Figure 2 the impact of PBH on 21-cm signal is primarily driven by heating. From the middle panel we see that PBH heating increases with the PBH mass, especially for $M_{\mathrm{PBH,ini}}>10M_{\odot}$, and we find that PBH mass evolution can increase gas temperature by a factor of 7 relative to the case without accounting for mass evolution. It is also interesting to note that PBH radiation may even heat $T_{\rm K}$ to be above $T_{\rm CMB}$, potentially flipping the sign of $T_{\rm 21}$ and turn it into an emission signal ($T_{21}>0$). Upcoming measurements can test this in MWA [77], MeerKAT [78] and SKA [79], etc. The full PBH-included $T_{21}$ evolution during EoR requires a more involved simulation, which we will postpone to a follow-up study.
@@ 4 Results
Figure 3 shows our upper limits on $f_{\rm PBH,ini}$ for $M_{\rm PBH,ini}$ within $[1,10^{4}]M_{\odot}$ range. Our main result is shown in the light blue shaded region, in which we properly account for mass evolution. From our discussions in Sec. 3, it can be inferred that PBH contribution to ionization through deposition term $[{\rm d}E/{\rm d}V{\rm d}t]_{\rm dep,c}$ term is proportional to X-ray emissivity $\epsilon_{\rm X}$, and from Eq. (3.4) and Sec. 2.2 one can easily show that $\epsilon_{\rm X}$ follows the scaling relation,
@@
where $\beta$ equals 2.3 and 0.25 for $\dot{m}\leq 0.02$ and $\dot{m}\geq 0.02$, respectively. $M_{{\rm PBH,ini}}$ scaling of our $f_{{\rm PBH,ini}}$ constraints can be well explained by Eq. (4.3). At smaller $M_{{\rm PBH,ini}}$ when mass growth is negligible ($M_{\rm PBH}/M_{{\rm PBH,ini}}\sim 1$), $\epsilon_{\rm X}$ is dominated by $\dot{m}$ term which is found to increase with PBH mass, thus from the black solid line we see that for $1M_{\odot}\leq M_{{\rm PBH,ini}}<10M_{\odot}$, our $f_{{\rm PBH,ini}}$ limit tightens from $10^{-2}$ to $3\times 10^{-4}$. At higher masses, accretion increases both PBH mass ($M_{\rm PBH}/M_{{\rm PBH,ini}}\gg 1$) and accretion rate itself ($\dot{m}$), therefore $f_{{\rm PBH,ini}}$ limit further tightens until $M_{{\rm PBH,ini}}\sim 10^{3}M_{\odot}$. At even higher masses ($M_{{\rm PBH,ini}}\gtrsim 10^{3}M_{\odot}$), accretion rate $\dot{m}$ starts to approach the Eddington limit $\dot{m}=10$. For the most extreme case in which $\dot{m}$ consistently takes $\dot{m}=10$, $M_{\rm PBH,ini}/M_{{\rm PBH,ini}}$ and $\dot{m}$ terms in Eq. (4.3) no longer depends on $M_{{\rm PBH,ini}}$ (see also Eq. (2.22)), thus we see that our constraints gradually flatten at $M_{{\rm PBH,ini}}\gtrsim 10^{3}M_{\odot}$, reaching $f_{{\rm PBH,ini}}\leq 2\times 10^{-7}$ at $M_{{\rm PBH,ini}}=10^{4}M_{\odot}$.
In the dark blue shaded region of Figure 3, we also show constraints without accounting for mass evolution. For this case Eq. (4.3) reduces to $\epsilon\propto\dot{m}^{\beta}$, since massive PBHs exhibits higher accretion rate $\dot{m}$, our constraint tightens for higher $M_{{\rm PBH,ini}}$ until it gradually flattens when $\dot{m}$ starts to approach the constant upper limit 10. At $M_{\odot}\leq M_{{\rm PBH,ini}}<2M_{\odot}$ the relevant $f_{{\rm PBH,ini}}$ limit is identical to that of with evolution. At higher masses, mass evolution boosts X-ray emissivity, therefore we see that when ignoring mass growth, $f_{{\rm PBH,ini}}$ constraints can be weakened by up to one order of magnitude.